A certain medical test is known to detect 51% of the people who are afflicted with the disease Y. If 10 people with the disease are administered the test, what is the probability that the test will show that:

All 10 have the disease, rounded to four decimal places?

At least 8 have the disease, rounded to four decimal places?

At most 4 have the disease, rounded to four decimal places?
Answer #1

X ~ bin ( n , p)

Where n = 10 , p = 0.51

binomial probability distribution is

P(X) = nCx px ( 1 - p)n-x

a)

P(X = 10) = 10C10 * 0.5110 * ( 1 - 0.51)0

= 0.0012

b)

P(X >= 8) = P(X = 8) + P(X = 9) + P(X = 10)

= 10C8 * 0.518 * ( 1 - 0.51)2 +10C9 * 0.519 * ( 1 - 0.51)1 +10C10 * 0.5110 * ( 1 - 0.51)0

= 0.0621

c)

P(X <= 4) = P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) + P(X = 4)

= 10C0 * 0.510 * ( 1 - 0.51)10 +10C1 * 0.511 * ( 1 - 0.51)9 +10C2 * 0.512 * ( 1 - 0.51)8

+10C3 * 0.513 * ( 1 - 0.51)7 +10C4 * 0.514 * ( 1 - 0.51)6

= 0.3526